find all solutions between 0 2pi for 7 cos(2x-3)=2
Answer by lwsshak3(11628)

find all solutions between 0 2pi for
7 cos(2x-3)=2
cos(2x-3)=2/7
using calculator set to radians
2x-3≈1.28 and 5.00 (in quadrants I and IV where cos>0)
2x≈4.28
x≈2.14
or
2x-3≈5.00
2x≈8.00
x≈4.00